OpenStudy (anonymous):

Jim has received scores of 74and 80 on his first two 100 point tests. What score must he get on his third 100 point test to keep an average of 83 or greater? Third test >= 'blank' (If the average is greater then the possible number of points, enter DNE)

OpenStudy (anonymous):

Ok, so (74 + 80 + x)/3 = 83 - does that make sense?

myininaya (myininaya):

inequality

OpenStudy (anonymous):

So do I divide the three then?

OpenStudy (anonymous):

You're right myin, but since we know the limit, 100 - aren't we just finding the minimum number that will give you an 83 average? Or am I reading/doing this wrong?

OpenStudy (anonymous):

Lulu, you multiply both sides of the equation by 3

OpenStudy (anonymous):

(Which gets rid of the /3 portion on the left)

OpenStudy (anonymous):

the answer doesn't have to be 83, it can be a higher average

myininaya (myininaya):

we want the average to be greater than or equal to 83 so we have \[\frac{74+80+x}{3} \text { is the average } \] we want \[\frac{74+80+x}{3} \text{ we want this to be greater than or equal to 83} \] symbolically this looks like \[\frac{74+80+x}{3} \ge 83 \]

myininaya (myininaya):

but once you get x=100 you could have also wrote x>=100 in the end so your way would have worked fine i just like my equations/inequalities read like my sentences

myininaya (myininaya):

that was an example by the way with the x=100

myininaya (myininaya):

i didn't solve this sorry lol

OpenStudy (anonymous):

So, I have a question that might be a bit silly. How do I work with the division of the equation. For the life of me I can't remember how to work an equation like that.

OpenStudy (anonymous):

Hey lulu - sorry I stepped away for a bit - so your question is regarding how to deal with the division of the equation?

OpenStudy (anonymous):

Yes, I'm not sure how to do it. Haven't covered it in a while and I can't remember how to go about it.

OpenStudy (anonymous):

No problem! So, the rule of thumb is when you multiply/divide, you need to do the same thing to both sides of the equation.

OpenStudy (anonymous):

So, in our case, you want to multiply both sides of the equation by 3, like this: \[(74 + 80 + x)/3 *3 = 83 * 3\]

OpenStudy (anonymous):

Oooh I get it, thank you so much!

OpenStudy (anonymous):

Awesome - thanks for sticking with me, I got confused myself there for a bit :P. And sorry for going quiet for a bit there - great work lulu, thanks!